OpenStudy (anonymous):

Simplify (a + 1)!/(a - 2)!

DivineSolar (divinesolar):

This is Prime, so it cannot be simplified any further.

DivineSolar (divinesolar):

See how?

OpenStudy (anonymous):

There's an answer in the back of the book though...

DivineSolar (divinesolar):

What are the choices?

OpenStudy (anonymous):

I just don't know how to get there It's a(a+1)(a-1) but I don't understand how to get that

DivineSolar (divinesolar):

a(a+1)(a-1) is what we are solving then?

OpenStudy (anonymous):

No, the question is: Simplify (a+1)!/(a-2)! And the answer that the back of the book gives is: a(a+1)(a-1) And I don't understand how to get that

OpenStudy (zarkon):

\[(a+1)!=(a+1)\cdot a\cdot (a-1)\cdot (a-2)!\]

OpenStudy (anonymous):

I'm still confused...

OpenStudy (zarkon):

do you understand what I typed?

OpenStudy (anonymous):

Yes, I think so

OpenStudy (zarkon):

\[\frac{(a+1)!}{(a-2)!}=\frac{(a+1)\cdot a\cdot (a-1)\cdot (a-2)!}{(a-2)!}\]

OpenStudy (zarkon):

\[=\frac{(a+1)\cdot a\cdot (a-1)\cdot \cancel{(a-2)!}}{\cancel{(a-2)!}}\] \[=(a+1)\cdot a\cdot (a-1)\]

OpenStudy (anonymous):

Oh! I got it! Thank you so much :)

OpenStudy (zarkon):

np